OpenStudy (anonymous):

Find the slope and y-intercept of the line x+y=5. the slope of the line is: _________ the y-intercept of the line is: _________ Please explain how to solve!

OpenStudy (anonymous):

to find the Yintercept set X to be 0 so 0 + y = 5 so ur Yinter is 5

OpenStudy (anonymous):

for the slope move the x value to the other side x+y=5 y=x+5 if u know the standar form of a liner eqution is... y=mx + B m= slove so ur slope in this case is just 1 since is the number infront of x now tht the equation is in standar form

OpenStudy (anonymous):

Thank you once again!!

OpenStudy (anonymous):

d u hve any questions thou?